A capillary tube of 0.4 mm diameter is placed vertically inside (i) water of surface tension 6.5 x 10-2 Nm-1 and zero angle of contact, (ii) a liquid of density 800 kgm-3, surface tension 5.0 x 10-2 Nm-1 and an angle of contact 30 degrees. Calculate the height to which the liquid rises in the capillary in each case.
1
Expert's answer
2020-10-23T09:09:14-0400


Let "r" be the radius of the capillary tube, "h" the rise of the liquid in the capillary, "\theta" the angle of contact.

The liquid makes contact with the tube along a line of length "2\pi r" . As the surface tension "S" acts tangentially to the liquid surface, its vertical component is "S \cos\theta" so that the total upward force due to it is "2\pi r S\cos\theta" . This balances the weight of a liquid cylinder of height "h" and weight "\pi r^2h\rho g" , where "\rho" is the density of the liquid.

"\therefore 2\pi rS\cos\theta=\pi r^2h\rho g\\
\Rightarrow h=\frac{2S\cos\theta}{r\rho g}\qquad ...........(1)"

Diameter of the capillary tube = 0.4 mm

"\therefore" Radius of the capillary tube, "r=0.2\ mm = 0.2\times 10^{-3}\ m"

(i) For water:

Surface tension of water, "S=6.5\times 10^{-2}\ N/m"

Contact angle, "\theta=0\degree"

Density of water, "\rho = 1000\ kg/m^3"

Substituting the values in Eq.(1),

"h=\frac{2\times 6.5\times 10^{-2}\times \cos0\degree}{0.2\times 10^{-3}\times 1000\times 9.8}\ m\\
\Rightarrow h=6.6\times 10^{-2}\ m\\
\Rightarrow h=6.6\ cm"

(ii) For the liquid:

Surface tension, "S=5.0\times 10^{-2}N/m"

Contact angle, "\theta=30\degree"

Density of the liquid, "\rho=800\ kg/m^3"

"\therefore h=\frac{2\times 5.0\times 10^{-2}\times \cos30\degree}{0.2\times 10^{-3}\times 800\times 9.8}\ m\\
\Rightarrow h=5.5\times 10^{-2}\ m\\
\Rightarrow h=5.5\ cm"

Answer: (i) Water rises to 6.6 cm in the capillary and (ii) the liquid rises to 5.5 cm